# Generalizing in Net-Zero Microgrids: A Study with Federated PPO and TRPO

### v1

@@ Abstract
This work addresses the challenge of optimal energy management in microgrids through a collaborative and privacy-preserving framework. We propose the FedTRPO methodology, which integrates Federated Learning (FL) and Trust Region Policy Optimization (TRPO) to manage distributed energy resources (DERs) efficiently. Using a customized version of the CityLearn environment and synthetically generated data, we simulate designed net-zero energy scenarios for microgrids composed of multiple buildings. Our approach emphasizes reducing energy costs and carbon emissions while ensuring privacy. Experimental results demonstrate that FedTRPO is comparable with state-of-the-art federated RL methodologies without hyperparameter tunning. The proposed framework highlights the feasibility of collaborative learning for achieving optimal control policies in energy systems, advancing the goals of sustainable and efficient smart grids.
keywords:
smart grids, federated learning, reinforcement learning, energy efficiency
journal:
[…]
{Frontmatter}
[…]
Cuadrado et al.
[…]
The concept of smart grids involves a series of strategies that encompass new dynamics that the electricity grid should integrate to adapt to the requirements introduced by several factors: climate change, increased integration on renewables, changes in regulations, and the need to improve energy efficiency, among others Ton and Smith (2012). Each component of that vision brings challenges we need to solve to make it feasible, and we are tackling one of them in the present work: optimal control of energy assets in microgrids, which is essential to achieve zero energy waste and to guarantee fast integration of renewables without affecting the reliability of the supply. This research impacts Sustainable Development Goals 7 and 11.
@@ 1 Introduction
Year after year, the majority of countries renew their commitment to the Sustainable Development Goals (SDGs) and the Paris Agreement in 2015 by evaluating the advancement in the established ambitious targets to combat climate change, with the eye set on the objective of limiting global warming to below 2°C, ideally 1.5°C, compared to pre-industrial levels. However, the most recent report from the Intergovernmental Panel on Climate Change (IPCC) IPCC (2023) reveals that humanity’s efforts remain insufficient, with many targets far from being achieved, which has become a constant. Many challenges need urgent action, and we should increase our efforts to get back on track. Machine learning (ML) is a technology that, when used correctly, can accelerate the actions needed to fight global warming. ML can support the transition from fossil fuels, accelerate renewable energy adoption, improve energy efficiency, and promote sustainable transportation, offering powerful tools to meet some of the most pressing climate goals Rolnick et al. (2019).
Integrating renewable energy sources like wind and solar into the energy grid is a significant challenge due to their intermittent and unpredictable nature, complicating real-time grid balancing. Fluctuations in renewable energy production can lead to supply-demand mismatches, necessitating advanced solutions such as energy storage systems and predictive analytics Perera et al. (2020); Gernaat et al. (2021). Smart Grid technologies, which leverage data-driven approaches, offer a way to address these complexities. By monitoring and managing the energy supply chain from generation to consumption, Smart Grids aim to optimize operations for diverse stakeholders, minimize costs, reduce greenhouse gas emissions, and ensure grid stability, reliability, and resilience.
Artificial Intelligence (AI) has gained recognition as a powerful tool for addressing the urgent threats of climate change. Its ability to process and analyze vast datasets allows for better comprehending environmental shifts and developing strategic solutions. Reinforcement Learning (RL), a subset of AI, offers some capabilities that are especially relevant to addressing the complexities of energy systems. Namely, its ability to handle non-linear, dynamic, and uncertain systems makes it particularly effective for managing distributed energy resources and microgrids. Model-free RL algorithms learn from interactions between an agent and its environment, enabling them to identify optimal control actions even under dynamic environments with constantly changing conditions without depending on a precise model, which is the traditional approach.
However, RL also faces challenges that hinder its applicability in real life. It is a method that relies heavily on massive sampling, making a data-hungry approach; as a consequence of its sample inefficiency, its performance decays when used in out-of-the-distribution (OOD) data (generalization) and particularly for applications in the energy efficiency field, RL needs to access sensitive data on energy usage to reach optimal solutions, creating the need of implementing privacy-preserving methodologies with a minor impact on the utility.
This work studies the application of RL within microgrids, which is one of the strategies that belong to the smart grid concept. The Department of Energy of the United States provided a commonly accepted definition for microgrid: “a group of interconnected loads and distributed energy resources within clearly defined electrical boundaries that act as a single controllable entity concerning the grid. A microgrid can connect and disconnect from the grid to operate in grid-connected or island mode” Ton and Smith (2012). For this purpose, we defined a custom microgrid scenario using a well-known environment called CityLearn Vázquez-Canteli et al. (2019), which allowed us to study the properties and challenges of RL and how collaborative learning can improve the learning process.
We propose a method based on two commonly used methodologies in the ML domain: RL and federated learning (FL). The first follows the intuition that model-free learning is a feasible alternative to managing distributed energy resources (DERs) in a microgrid. More specifically, we picked the Trust Region Policy Optimization (TRPO) Schulman et al. (2015) algorithm to enjoy the benefits of its adaptive step size, considering we are not using large models. The second is to profit from having multiple agents learn in parallel while preserving their privacy. This framework, which we named FedTRPO, harnesses inherent generalization capability and privacy guarantees from FL to discern latent relationships within the feature space. We present a summary of our contributions next:
@@
  • We present a customization to the CityLearn environment that allows a better understanding of the learned policies and the impact of the reward function in this case. We also modified some characteristics that better represent the dynamics of real-life scenarios.
  • This work introduces the FedTRPO learning framework and evaluates it in the context of our proposed smart grid problem, leveraging the faster convergence of TRPO with the fewer communication rounds time of FL.
@@
  • Our results demonstrate the generalization capabilities of FL and how reasonable considerations like personalization and split-learning Han et al. (2022) can reduce the optimality gap in the learned policies.
  • We compared our framework against state-of-the-art RL methodologies in a standard FL setup.
@@
  • We provide a wrapper for the CityLearn environment that makes it compatible with TorchRL Bou et al. (2023), improving the modularity and portability of our code. We also provided a version of the TRPO algorithm using TorchRL.
@@ 2 Related Work
Federated Learning. FL provides a collaborative learning framework relying on clients that share the same model architecture but learn locally based on their portion of the data distribution. After some training steps, reducing the clients’ models generates a new iteration of the original model that combines the individual learnings. Federated Averaging (FedAvg) McMahan et al. (2017) stands as the pioneering framework in FL, which averages the weights using a centralized server, often called the parameter server (PS). Other approaches offer a decentralized way of averaging the models; for instance, the authors of Koloskova et al. (2019) introduced a methodology that consists of direct communication between clients for several rounds until they all update their model with the reduced one, which they called Gossip algorithms. Further research on FL tackles some of the challenges of FedAvg, like the suboptimal nature of averaging models under heterogeneity of the clients (difference in the portion of the data distribution they access) or the privacy loopholes created by FedAvg, contrary to the belief that it preserves the privacy of the clients. The work in Vepakomma et al. (2018) presents an interesting approach named split-learning, which tackles personalization (adjusting the combined model to the client’s characteristics) and privacy by keeping a local private part of the model not shared for aggregation. Differential Privacy (DP) Geyer et al. (2017) is another technique often used with FL. It consists of offuscating the original data by passing it through a noise mechanism that ensures a level of utility while ensuring the privacy preservation of individuals or groups to a certain extent. Other topics of interest in the domain are learning in heterogeneous setups Diao et al. (2021) and analysis of the generalization capabilities of FL Yuan et al. (2022). With the proper considerations, FL can handle heterogeneous data training, preserve privacy, and generalize on unseen data, making it a promising methodology for addressing challenges related to smart grid scenarios.
RL-based and Federated RL-based microgrids. The microgrid scenario increasingly receives attention in research on the intersection of ML and energy. The most common setup consists of a group of homes constituting a microgrid. Each one of them has its load profile (which could be controllable or not), energy generation resources (generally solar panels or wind turbines), and storage devices (electricity, heating, or cooling.) There are community efforts that enable tackling this scenario, and the CityLearn environment Vázquez-Canteli et al. (2019); Nweye et al. (2024) is the most recurrently used. It provides an interface compatible with Farama Foundation’s Gymnasium Towers et al. for building energy coordination, vehicle-to-grid control, thermal control, and demand response based on a real dataset Wilson et al. (2021). Several contributions used it in the research of RL; for instance, the work presented in Nweye et al. (2023) explored a multi-agent RL approach for training agents to control the battery energy storage systems (BESS) pursuing the minimization of a weighted sum of a set of metrics of interest for energy user and grid operator: electricity price, carbon emissions, zero net energy, average daily peak, and load factor. To do so, they implemented a multi-agent version of the Soft Actor-Critic algorithm with a shared replay buffer. They employed three different setups for the learning case, differing in the number of agents contributing to the memory buffer and the data points used for the training. The authors of Nweye et al. (2022a) explored the challenges of multi-agent RL using the same environment, showing that a warm start with rule-based control (RBC) policies seems to be the best way of directing the RL learning to find feasible RL policies. However, other contributions leveraged their environments using private datasets of energy usage, like the work presented by Ceusters et al. (2021), which adopted a model-based reinforcement learning (RL) approach tailored to multi-energy systems, enhancing it with constrained optimization algorithms to derive optimal solutions. Meanwhile, Su et al. (2022) introduced a Federated RL (FRL) methodology, considering the dynamics between Energy Data Owners (EDOs) and Energy Service Providers (ESPs), addressing free-rider phenomena and non-independent and identically distributed (IID) data distributions. Furthermore, Cuadrado et al. (2023a) introduced a hierarchical FRL, which demonstrated advantages in the learning process while integrating the FedAvg method in an existing hierarchical RL setup detailed in Cuadrado et al. (2023b).
Although we consider that the existing work has been paving the foundations of RL applicability in real life, we find two relevant concerns. First, many contributions aim to see a monotonic increase behavior of the return function (the main objective of RL) to claim that a policy is optimal when that is not necessarily the case. One of the main challenges of this scenario is avoiding the local minima from trivial cases (e.g., not doing anything, always charging, constantly discharging, on-off policies), and an increased return could not mean optimality, not even gradients of magnitude close to zero mean optimality. Second, we noticed that in the contests organized by the same authors of CityLearn in the last years, the winning teams disregarded the impact of the emissions, prioritizing other metrics that had higher weight in the final score function Nweye et al. (2022b); Nagy et al. (2021). We found that in the context of climate change, the reduction of emissions should be a priority, and we might be ignoring why this scenario seems harder to resolve when prioritizing emission reduction.
@@ 3.1 Environment
As mentioned before, we used CityLearn environment to try our framework, however, instead of using the standard data they include in their package we decided to generate synthetic data that allow us to understand clearly the learned policies through simple cases. For this work we focused on data representing homes that can be self-sustainable, in other words, each home can provide enough energy to serve its demand, so the optimal solution must have zero cost and emissions. Our data generation process takes as a reference the citylearn_challenge_2022_phase_all scenario. It modifies the time series for non-shiftable load to ensure solar energy can cover the demand profile completely. We included other modifications to the environment:
  • We added a new logic for the selling energy rate. We found it unrealistic that in the original scenario, the users receive a compensation equal to the Time-of-Use (ToU) tariff. This setup discourages houses from being self-sustainable, as we will detail later.
  • We modified the capacity of the BESS devices. We simplified their models so they don’t suffer degradation or decay in their performance, helping to ease the understanding of the policies we are training, which is one of the main interests of this research.
  • We introduced additional observations representing load predictions for some steps ahead, which we consider essential to finding optimal policies.
@@
We kept the grid price pattern, which fluctuates throughout the day, peaking during high-demand hours, while the solar panels contribute energy during daylight hours. Allowing the agent to stockpile energy during off-peak times and, when solar generation exceeds immediate needs, later release stored energy during high-price periods. Emissions follow the actual emission rate from the grid mix (kgCO2e/kWh) for the dataset used in the environment. With our process, we can generate training data up to the number of buildings existing data in the reference case. We generate convex combinations of the training buildings’ time series for the validation data. Each episode spans twenty-four hours, and $t$ is the index for the time step. For a building $i$, an agent considers the accessible observations to decide the charge or discharge of the battery in a time step ($E^{\text{batt}}_{t,i}$), aiming to meet the building’s load requirements. Its primary objective is minimizing its house’s overall energy cost and emissions and, as an extension, for the whole microgrid. The table 3.1 presents an overview of the RL case.
[…]

### v2

# Generalising Battery Control in Net-Zero Buildings via Personalised Federated RL

@@ Abstract
This work studies the challenge of optimal energy management in building-based microgrids through a collaborative and privacy-preserving framework. We evaluated two common RL algorithms (PPO and TRPO) in different collaborative setups to manage distributed energy resources (DERs) efficiently. Using a customized version of the CityLearn environment and synthetically generated data, we simulate and design net-zero energy scenarios for microgrids composed of multiple buildings. Our approach emphasizes reducing energy costs and carbon emissions while ensuring privacy. Experimental results demonstrate that Federated TRPO is comparable with state-of-the-art federated RL methodologies without hyperparameter tuning. The proposed framework highlights the feasibility of collaborative learning for achieving optimal control policies in energy systems, advancing the goals of sustainable and efficient smart grids. Our code is accessible in this repo.
Net-zero, Federated Learning, Personalization, Reinforcement Learning
@@ 1 Introduction
Global reviews of progress under the SDGs and the 2015 Paris Agreement still show the world drifting beyond the 1.5–2 °C ceiling, as confirmed by the latest IPCC assessment (IPCC, 2023). Because building operations consume about 30% of global final energy and produce roughly 26% of energy-related CO2 emissions (International Energy Agency, 2024), cutting their footprint is pivotal. A recent analysis estimates that AI-enabled controls alone could trim building energy demand and emissions by 8–19% by 2050 (Ding et al., 2024). Machine learning (ML) is a technology that, when used correctly, can accelerate the actions needed to fight global warming. ML can support the transition from fossil fuels, accelerate renewable energy adoption, improve energy efficiency, and promote sustainable transportation, offering powerful tools to meet some of the most pressing climate goals (Rolnick et al., 2019).
Rooftop solar and other on-site renewables introduce fast power swings that can create acute shortage events at the building level (Zheng et al., 2024; Perera et al., 2020; Gernaat et al., 2021). Modern Building Energy Management Systems (BEMS) mitigate these fluctuations by pairing local storage with predictive control and have been shown to cut energy use and costs across diverse facilities (Michailidis et al., 2024). Model-free reinforcement learning (RL) can learn such control policies directly from data, capturing the non-linear dynamics of distributed resources without a detailed physics model (Barbalho et al., 2025). Yet standard RL is sample-hungry and generalises poorly to unseen climates or occupancy patterns (Bekal et al., 2025). Privacy-preserving federated approaches overcome both issues by exchanging model updates rather than raw traces, protecting occupant data while pooling experience across sites (Lee et al., 2021). Recent demonstrations report cost and CO2 reductions when scaling federated RL to dozens of heterogeneous buildings in multiple climates (Sievers et al., 2025a). Privacy guarantees remove regulatory and organisational barriers to pooling knowledge across many buildings, making collaborative learning a practical path to scalable net-zero control (Sievers et al., 2025b).
This study explores building-scale net-zero microgrids (“a group of interconnected loads and distributed energy resources within clearly defined electrical boundaries that acts as a single controllable entity” (Ton & Smith, 2012)) as test-beds for scalable, privacy-preserving control. We generated scenarios where buildings are self-sustainable by design as long as the electrical storage is managed using the optimal policy. Using a custom microgrid scenario based on CityLearn (Vázquez-Canteli et al., 2019), we explored the generalization capabilities of RL agents and the challenges they faced even in simplified scenarios showing that naive policies are difficult to learn in a collaborative way. We improved the learning process using Federated Learning (FL), in this setup each building trains a local RL agent but only shares parameter updates, a practice that retains data privacy while exploiting the scalability and generalisation benefits identified in recent federated-learning surveys (Kairouz et al., 2021; Lee et al., 2021). Additionally, we included personalisation through split-learning (Han et al., 2022), further shrinking the generalisation gap that often appears when controllers are deployed in heterogeneous conditions or occupancy patterns (Bekal et al., 2025). We present a summary of our contributions next:
@@
  • We introduce a custom environment based on CityLearn, which allows a better understanding of the learned policies and the impact of the reward function in the learning process. We also modified some characteristics that better represent the real-life dynamics.
  • We provide experimental assessment of traditional RL algorithms in our environment, testing generalization and optimality of the learned policies.
@@
  • Our results demonstrate the generalization capabilities of FL and how reasonable considerations like personalization and split-learning (Han et al., 2022) can reduce the optimality gap in the learned policies.
@@
  • We provide a wrapper for the CityLearn environment that makes it compatible with TorchRL (Bou et al., 2023), improving the modularity and portability of our code. We also provided a version of the TRPO algorithm using TorchRL.
@@ 2 Related Work
Federated Learning. FL trains a common model by letting each client update locally and then merging weights. The seminal FedAvg algorithm performs this aggregation on a central server (McMahan et al., 2017), whereas Gossip protocols average peer-to-peer without a server (Koloskova et al., 2019). Subsequent work addresses FedAvg’s limits under non-IID data and privacy leakage: split learning keeps part of each network private to enable personalisation (Vepakomma et al., 2018), and differential privacy injects calibrated noise into updates (Geyer et al., 2017). Studies on heterogeneous clients (Diao et al., 2021) and generalisation theory (Yuan et al., 2022) confirm that, with these extensions, FL can learn robust, privacy-aware controllers—making it well suited to smart-grid scenarios with diverse buildings.
RL-based and Federated RL-based microgrids. Recent studies model microgrids as groups of houses with distinct loads, rooftop renewables, and storage devices, typically using the open-source CityLearn platform (Vázquez-Canteli et al., 2019; Nweye et al., 2024) built on Gymnasium (Towers et al., ) and real demand data (Wilson et al., 2021). Within CityLearn, (Nweye et al., 2023) trained multi-agent Soft Actor-Critic controllers for battery systems to cut cost, carbon, and peak demand, while (Nweye et al., 2022a) showed that warm-starting RL with rule-based policies speeds convergence. Beyond public benchmarks, a model-based RL approach with constrained optimisation targets multi-energy hubs (Ceusters et al., 2021); a federated RL scheme addresses free-riders and non-IID data (Su et al., 2022); and hierarchical FRL embeds FedAvg in layered control architectures (Cuadrado et al., 2023a, b). Collectively, these works highlight both the promise and the outstanding challenges of RL and FRL for collaborative, net-zero microgrids.
Despite steady progress, two gaps remain. First, many studies treat a monotonic rise in cumulative reward as proof of optimality, yet trivial local minima, such as always charging or doing nothing, can produce similar curves. Second, CityLearn challenge winners (2021–2023) optimised the highest-weighted terms and largely ignored CO2 impact (Nweye et al., 2022b; Nagy et al., 2021), even though emission reduction should be central in climate-aligned control.
@@ 3.1 Environment
We used our customized version of the CityLearn environment. We decided to generate synthetic data that allows us to understand clearly the learned policies through simple cases, focusing on representing homes that can be self-sustainable, in other words, each home can provide enough energy to serve its demand, so the optimal solution must have zero cost and emissions. Our data generation process takes as a reference the citylearn_challenge_2022_phase_all scenario. It modifies the time series for non-shiftable load to ensure solar energy can cover the demand profile completely. We included other modifications to the environment:
  • New logic for the selling energy rate. We found it unrealistic that in the original scenario, the users receive compensation equal to the Time-of-Use (ToU) tariff. This setup discourages houses from being self-sustainable.
  • We simplified the storage models so they don’t suffer degradation or decay in their performance, helping to ease the understanding of the policies we are learning, which is one of the main interests of this research.
  • We introduced additional observations representing load predictions for some steps ahead.
We kept the grid price pattern in the original CityLearn environment, which fluctuates throughout the day, peaking during high-demand hours, while the solar panels contribute energy during daylight hours. Allowing the agent to stockpile energy during off-peak times and, when solar generation exceeds immediate needs, later release stored energy during high-price periods. Emissions follow the actual emission rate from the grid mix (kgCO2e/kWh) following the original dataset used in CityLearn. We generate convex combinations of the training buildings’ time series for the validation data. Each episode spans twenty-four hours, and $t$ is the index for the time step. For a building $i$, an agent considers the accessible observations to decide the charge or discharge of the battery in a time step ($E^{\text{batt}}_{t,i}$), aiming to meet the building’s load requirements. Its primary objective is minimizing its house’s overall energy cost and emissions and, as an extension, for the whole microgrid. The table B in the Appendix presents an overview of the RL case. Our design of the reward function is a weighted sum of the resulting cost and emissions for the complete episode (a full day), and we established the values as $W_{C}=0.4$ and $W_{G}=0.6$ to enforce the prioritization of emission reduction. We also added a penalization to the reward function to guide the agent to avoid unfeasible actions.
3.2 Model
We explore the advantages of a collaborative learning approach for the case of microgrids, given their distributed nature: each household has an agent that explores and learns from different sections of the state space, given the individual attributes of the household (generation, load profile, battery capacity). We have two types of models for both Policy and Critic networks: a shallow neural network and a custom model that adds a grouping feature logic as the first part of the feature extraction; we refer to it as Personal Encoding in Figure 1. This block represents a part of the network that remains local; the FedAvg process does not access it. It encodes features that should remain private (e.g., number of people living in the household, ToU strategy, appliances, area of the house, etc.) and learns a latent representation, used as an input to the public model, stacked with the public features, which aligns with the concept of split-learning (Vepakomma et al., 2018). We can assume it as a personalization strategy.
4 Experiments
@@
[…]
4.1 Experiment Setup
We generated two environments for our experiments using the simplified version of the data 3(a): a microgrid with two houses and one with five houses. Each one also has a shifted version that creates day gaps between the house’s time series, simulating that each house is subject to different environmental conditions and solar irradiance patterns. Regarding the model configurations, as described 1, we run experiments in four setups: shallow network, shallow network with personal encoding, shallow network with grouping, and shallow network with personal encoding and grouping. We defined four group features: load-related, solar-related, pricing-related, and the rest. We fixed the communication rounds for both algorithms we tried, Proximal Policy Optimisation(PPO) and Trust Region Policy Optimisation (TRPO), to evaluate how close they can come to the optimal solution within the same number of iterations. Another essential detail is setting the discount factor $\gamma=1$, meaning that our agents don’t consider a discounted reward, given that we are running episodes of just one day (24 time steps).
4.2 Quantitative Results
The tables 2 and 2 show the experiments we mentioned in the previous section: base means we used the shallow network, pe that we included personal encoding, and gf that we grouped features. We ran each combination for five different seeds. We averaged the results, making $160$ experiments with an average run time of $70m$ for the PPO runs, and $110m$ for the TRPO runs using an NVIDIA RTX 4000 Ada Generation GPU.
In the five-building environment, grouping features seems to play a more critical role than in the case of just having two buildings. We interpret this as the effect of having more agents, which means more sampling and more diverse data, which can help the networks unveil patterns in the data that are not evident by simple inspection. However, this does not seem to be the case for TRPO, in which the simple model yields the best performance, although it is essential to note that the performance of TRPO is poor based on the reward function. PPO overperforms TRPO significantly, but we highlight that we didn’t spend time fine-tuning it. Still, it managed to identify some patterns in the data, as we will show in the qualitative results in the Appendix.
5 Conclusion
We examined battery control in net-zero microgrids. Using a stripped-down CityLearn scenario with a known optimal policy, we studied RL combined with FedAvg and split-learning personalisation to balance privacy, heterogeneity, and policy quality. Federated TRPO performed on par with a tuned federated PPO: PPO was sometimes superior but stalled near the optimum, whereas TRPO converged reliably without hyperparameter tuning. A pragmatic remedy is to warm-start with a few PPO iterations and then switch to TRPO, leveraging TRPO’s adaptive step size in the final optimisation phase. The study shows that even “simple” optimal policies are surprisingly hard to learn, supporting curriculum designs that start with well-controlled cases before moving to full real-world complexity.
Future Work. Next steps include benchmarking against model-predictive control (MPC) and additional RL baselines to gauge relative sample-efficiency and robustness. We will design new “toy” microgrid cases—varying storage sizes, demand shapes, and renewable mixes—to probe policy behaviour and measure the optimality gap more precisely. Finally, we plan to scale experiments from five to dozens of buildings, testing how privacy, communication cost, and convergence behave in larger federations.
Appendix A Definition of RL Scenario
A.1 Reinforcement Learning
The RL component of our methods happens at the local training. We compared TRPO against a state-of-the-art algorithm called Proximal Policy Optimization (PPO) (Schulman et al., 2017), which emerged as an alternative to TRPO that avoids the use of second-order information. PPO relies on the same theoretical base as TRPO but redefines the policy optimization objective as an unconstrained expression. The scenario we study in this work, and the case of microgrids’ agents, does not require the use of very deep neural networks, and that is a reason why we believe that second-order methods are feasible in these cases, especially for RL methods, which tend to be very sensitive to the learning rate. TRPO explores the vast space by identifying suitable step sizes within a confined region, ensuring stable policy, monotonic improvement, and more effective exploration than other policy gradient methods.
Another essential aspect of opting for an RL approach is defining a proper reward function, especially considering that we defined a continuous action space, which means the number of possible actions is unbounded. The typical approach for this case is to use a policy network that outputs the parameters of a Gaussian distribution (reparameterization trick) to sample a real value that defines the action from that distribution. Throughout the training process, we don’t have guarantees that the policy network will explore feasible actions (within the range specified in B), especially in the first iterations. That fact motivated the addition of a penalization $P_{i,t}$ that impacts the return of a trajectory (the accumulated sum or rewards.) An essential consideration we made when designing our reward function is that we are not guiding it to follow a specific policy we know beforehand; we are just guiding the exploration of the agent by avoiding ineffective actions.
A.2 Federated RL
An actual FL setup implies having clients interact through a communication network channel throughout the training rounds. In the context of our study case, we can imagine it as the agents operating in an embedded device that controls the batteries of each household and connects to a parameter server through a dedicated network connection. However, this case would introduce some of the challenges of FL: limited bandwidth, faulty nodes or updates, Byzantine clients, etc. The main interest of this work is not to alleviate those challenges but to explore the benefits of collaborative learning. Thus, we are not required to implement the exact flow of FL to perform our experiments.
@@ A.2 Federated RL
Appendix B Synthetic Data
For this research, we decided to generate our data as it allowed us to explain better the policies we learned. Although using accurate data is always desirable for ML, there is low availability, especially regarding energy usage patterns, which concerns privacy. For example, the CityLearn environment constructs its 2022 scenario using data from 17 Sierra Crest Zero Net Energy community buildings from another study about ToU strategies (Narayanamurthy et al., 2016). They couldn’t access all the accurate data details, so they had to generate data that emulated the behavior based on some prior information about the ToU strategies the households might have used. Research about microgrid control systems is not particularly affected by lack of data, as prior information about consumption behaviors, typical weather conditions given a location, seasonal changes, and factors influencing energy use are accessible. Synthetic data could represent typical scenarios an agent may face and have the advantage of creating relevant evaluation scenarios by adding noise or making a convex combination of seen data.
(a) Simplified Dataset
(b) Normal Dataset
We generated our synthetic data by altering the Non-Shiftable Load time series using a simple rule that defines the policy we want to learn: the households have enough energy and battery capacity to serve their demand when there is solar energy and to store what is necessary when there is no sun. By definition, our scenario leads to net-zero energy consumption. Figure 3 shows how two versions of the data we generated, in the case of 3(a), we simplified the solar profile by making it constant through the sunny hours. For 3(b), we kept the solar profile as in the original reference case from CityLearn. By simplifying these profiles, we intended to reduce the state space for the policy while still keeping it challenging to learn. Another critical difference in our scenario is the pricing for energy. Our reference scenario from the CityLearn environment used the real pricing time series used in the Zero Crest Community Study. Still, for the environment implementation, they defined that the selling price is the same tariff whenever a household sends back energy to the grid. By inspecting 4(a), it is easy to notice that a policy optimizing exclusively by the energy cost would charge as much as possible in the low-price hours to sell energy when it is almost twice the buying price. In our scenario, we defined a margin of 40% of the lowest daily electricity pricing to ensure we don’t learn policies that bypass the weighting factor $W_{G}$ by finding profitable energy trades.
(a) Original scenario
(b) Our scenario